# Super-resolving frequency measurement with mode-selective quantum memory

### v1

@@ IV Methods
Figure 5: (a) Experimental setup. The apparatus consists of three main sections: pulse generation, memory interaction, and filtering and detection. Both signal and control pulses originate from a continuous wave (CW) laser at the control-field frequency. An electro-optic modulator (EOM), EOM 1, first carves the CW light into pulses. A 90:10 beam splitter (BS) then splits the laser into a high-power control path (90%) and a low-power signal path (10%). The laser in the signal path is later shifted to the signal frequency via EOM 2. The two beams are recombined via a beam displacer (BD) before entering the memory cell for storage and retrieval. Post-retrieval filtering comprises a Glan-Taylor polarizer (GTP), three double-passed etalons, and a band-pass filter (BP). The signal is finally routed to fiber beam-splitters and four superconducting nanowire single-photon detectors (SNSPDs) for detection. Employing multiple detectors helps mitigate detector dead time effects and enables higher overall count rates. (AWG: arbitrary waveform generator; PC: Pockels cell; PBS: polarizing beamsplitter; M: mirror; DM: dichroic mirror; HWP: half-wave plate; WPs: a half-wave plate and a quarter-wave plate.) (b) Pulse sequence. The sequence consists of three pulses: the first pulse is the signal input, which carries the spectral information to be estimated. The second pulse serves as the control read-in pulse, initiating the coherent storage process. The third pulse acts dually as both the control read-out pulse to retrieve the stored signal and a reference pulse for real-time signal power monitoring. The signal sequence is delayed by 437 ns to ensure temporal overlap of the signal input pulse with the control read-in pulse at the memory cell for storage. After a 200 ns storage period, the stored signal was retrieved using the control read-out pulse. During the measurement window, the pump laser was disabled, while the Pockels cell was activated exclusively during the retrieval window to suppress background from the pre-storage signal.
@@
Our Raman memory setup is shown in Fig. 5(a). The storage medium consists of a 75 mm-long cesium (Cs-133) vapor cell with a three-layer $\mu$-metal magnetic shielding. We heated the cell to $105\pm 1$°C using a quad-twisted cryogenic wire, achieving an effective optical depth of $(4.8\pm 0.1)\times 10^{3}$. The atoms were initialized to the initial state |$\ket{g}$ by an external cavity diode laser (ECDL, Toptica DL100) driving the D1 transition ($6^{2}\textbf{S}_{1/2}\rightarrow 6^{2}\textbf{P}_{1/2}$). An electro-optical modulator (EOM 3, EOSpace 850UL) gated the pump laser, switching it off during the experimental window. We achieved a pumping efficiency of $(99.6\pm 0.1)$%.
@@
We operated our memory with a detuning $\Delta=2\Delta_{\text{hf}}$ to suppress four-wave-mixing noise [55]. Both the signal and control fields were derived from a single continuous wave (CW) ECDL laser (Toptica DL Pro) operating at the control frequency (351.7122 THz). An arbitrary waveform generator (AWG, Tektronix AWG70001A) drove the electro-optical modulator EOM 1 (Sacher Lasertechnik AM830PF) to carve the CW laser into pulses. The laser was then split using a 90:10 beam splitter, directing 90% of the optical power to the control field path. The remaining 10% went to EOM 2 (New Focus 4851), which generated a sideband at the signal frequency (351.7030 THz). We selected this sideband as the signal field by using a Fabry-Pérot etalon with a free spectral range (FSR) of 36.8 GHz. The signal field was delayed by 430 ns using a long fiber to achieve temporal synchronization with the control pulses, ensuring their temporal overlap for memory storage and retrieval. Before entering the memory, the signal field also passed through a Pockels cell to remove residual CW background in the retrieval window. The polarization of the signal and control fields was set to orthogonal for memory interaction, allowing their combination using a beam displacer before being focused into the vapor cell.
We used a pair of convex lenses to focus the signal, control and pumping beams into the vapor cell to enhance the interaction strength. The beam widths at the focus were approximately $164\pm 5$ $\mu$m for the signal and $190\pm 5$ $\mu$m for the control. As shown in Fig. 5(b), the first signal pulse meets the second control pulse for storage. After a storage time of 200 ns, the second control pulse retrieves the stored signal. The third signal pulse serves as a reference for the signal pulse power. All signal and control pulses share the same Gaussian envelope width parameter as a fundamental Gaussian pulse (HG0) with an electric field full width at half maximum (FWHM) of 50 ns.
@@
After the memory, we suppressed the control and pump fields using a Glan-Taylor polarizer, followed by three double-passed etalons (two with FSR of 18.4 GHz and one with FSR 103 GHz) and a bandpass filter (central wavelength: 850 nm, FWHM Bandwidth: 10 nm). Finally, we split the signal into four equal parts and detected them using four superconducting nanowire single-photon detectors (SNSPDs, Photon Spot). The detected counts were registered using a time tagger (Swabian TimeTagger20).
@@
$\Psi(t)=A\cos\left(\frac{\epsilon\sigma t-\phi}{2}\right)\exp(-t^{2}\sigma^{2}),$
@@
One key requirement for this experiment is the accurate generation of signal and control pulses. To ensure high-fidelity pulse carving, we characterized the frequency response of our electronic pulse generation system, which comprises the AWG, an RF splitter, and two RF amplifiers that drive the EOM 1. By measuring the system’s frequency response function, we applied a correction to the input signals of the AWG, compensating for frequency-dependent variations in the RF components. This correction ensures uniform amplification across all frequency components, minimizing distortions in the carved pulses. The measured intensity pulse shapes of HG0 and HG1 are presented in the Supplementary Materials.
@@
For each signal, comprising one separation and one phase setting, we first stored it using a control pulse of HG0 with a pulse energy of $130\pm 1$ pJ. At this control energy, our memory typically operates with a storage efficiency of $(39.5\pm 0.6)\%$ and a total internal efficiency of $(14.6\pm 0.1)\%$ for a storage duration of 200 ns. The stored signal was retrieved using a control pulse in the HG1 mode, with the same pulse energy as the HG0 write-in pulse. The HG1 mode was chosen to minimize distortion in the pulse sequence due to its symmetrical temporal profile. Importantly, the choice of retrieval mode has a negligible impact on the retrieval efficiency. The pulse sequence was repeated every 3 $\mu$s, with a total measurement time of 2 s. To accurately account for signal background noise and dark counts, we also recorded data with the control pulse blocked for an equivalent duration. The counts detected in the retrieval time window during this control-blocked run were later used to subtract the noise counts from the retrieved signal counts to minimize the estimator bias caused by these noises. The above procedures were repeated for all four phases. Subsequently, we modified the control read-in pulse to the HG1 mode while keeping the read-out pulse unchanged and repeated the measurements. Finally, the entire experiment was repeated for all separations of interest.
@@ Acknowledgments
Supplementary Material: Super-resolving frequency measurement with mode-selective quantum memory
@@ IV.1 Fundamental bias at small separations
With the crosstalk matrix $\bm{M}$, the normalized measurement probability of HG1 projection is
$p_{1}=\beta-\frac{16(\alpha+\beta-1)}{16+\epsilon^{2}},$ (S1)
and the HG0 projection probability is $p_{0}=1-p_{1}$. The exact expression for the Fisher information of the non-ideal two-mode projection method can be derived from Eq. 2 in the main text as follows:
@@
We assume the mode crosstalk between the two modes is low, i.e. $\alpha\approx 1$ and $\beta\approx 1$, and the separation is small. Then Eq. S2 can be approximated as
$\mathcal{F}(\epsilon)\approx\frac{1}{4\left[(1-\alpha)/(\epsilon/4)^{2}+1\right]}.$ (S3)
Using maximum likelihood estimation (Eq. 5 in the main text), the likelihood function is given by
$\mathcal{L}=N_{0}\ln{p_{0}}+N_{1}\ln{p_{1}},$ (S4)
@@
Figure S1: Simulation results of the mean square error (MSE) and the unbiased Cramér-Rao lower bound (CRLB) for maximum likelihood estimation. (a) Simulated MSE, where the white region indicates the onset of truncation due to the non-negativity constraint. (b) MSE with the non-negativity constraint applied. (c) The unbiased CRLB. (d) The difference between the CRLB and the MSE (CRLB -- MSE).
Since we employ photon count detection, the detection counts $N_{0}$ and $N_{1}$ follow a Poisson distribution. Specifically, we model $N_{1}$ as $\text{Poisson}(\mu_{1})$, where $\mu_{1}=p_{1}N$. Substituting this distribution into the estimator formula in Eq. S5, we derive the mean square (MSE) error as
@@
The MSE simulated using our experimentally measured crosstalk is shown in Fig. S1 as a function of the separation $\epsilon$ and the number of photons $N$. In panel (a), the white region indicates where the MLE estimators begin to truncate negative estimates, and panel (b) displays the corresponding MSE with this truncation applied. Panel (c) presents the unbiased Cramér-Rao lower bound (CRLB). By taking the difference between the unbiased CRLB and the MSE (CRLB -- MSE), we obtain panel (d), which highlights a transition: initially, the MSE exceeds the unbiased CRLB (blue regions), but as the separation decreases and with larger bias, it drops below the CRLB (yellow regions).
@@ IV.2 Quantum memory interaction
[…]
@@
The modal properties of the Raman memory can be understood by performing singular value decomposition (SVD) on the Green functions $K_{1}$ and $K_{2}$ [56]. For the storage process, $K_{1}(z,\tau)$ can be decomposed into
@@ IV.4 Different retrieval modes and retrieval times
We also performed the experiment using different retrieval modes to demonstrate the flexibility in choosing the read-out control pulse. The results are shown in Fig. S4(a) and (b), where the control read-out pulses are set to HG0 and HG1, respectively. The superresolution performance remains comparable in both cases, and the retrieved signals (the second red pulse) clearly exhibit the corresponding temporal mode shapes.
@@ IV.5 Sources of mode crosstalk
IV.5.1 Memory storage efficiency
The primary source of temporal mode crosstalk arises from the memory’s storage interaction. Although the memory operates approximately in a single-mode regime, increasing the control field strength induces stronger AC Stark shifts. This leads to the coupling of orthogonal signal modes into the spin wave, distorting the ideal projection probabilities and introducing bias. Figure S5 illustrates how different storage efficiencies (i.e. different control read-in pulse power) affect the superresolution performance and estimation precision relative to DI methods.
Panel (a) shows the raw estimators $\hat{\epsilon}=4\sqrt{N_{1}/N_{0}}$ for different storage efficiencies. As storage efficiency increases, so does crosstalk, resulting in greater bias in the raw estimators. In panel (b), we apply maximum likelihood estimation and plot the corresponding unbiased CRLB for the two-mode demultiplexing method, alongside the quantum CRLB and the CRLB for DI methods. Higher storage efficiency leads to an increase in the CRLB, resulting in reduced estimation precision. This trend is further quantified in panel (c), where the precision improvement factor, $\mathcal{F}/\mathcal{F}_{\text{DI}}$, is plotted as a function of separation $\epsilon$. The mode demultiplexing method offers logarithmic improvements for small separations, but reaches an asymptotic limit as $\epsilon\to 0$. This limit defines the superresolution parameter [43], which is plotted against storage efficiency in panel (d). As storage efficiency decreases, the superresolution parameter increases logarithmically. However, very high precision improvement using very low storage efficiencies is impractical in experiments, as fewer signal photons are retrieved, constant background and control field leakage reduce the signal-to-noise ratio, potentially increasing the observed crosstalk and estimation bias.
IV.5.2 Control leakage

### v2

@@ IV Methods
Figure 5: Schematic of experimental setup. (A) Experimental setup. The apparatus consists of three main sections: pulse generation, memory interaction, and filtering and detection. Both signal and control pulses originate from a continuous wave (CW) laser at the control-field frequency. An electro-optic modulator (EOM), EOM 1, first carves the CW light into pulses. A 90:10 beam splitter (BS) then splits the laser into a high-power control path (90%) and a low-power signal path (10%). The laser in the signal path is later shifted to the signal frequency via EOM 2. The two beams are recombined via a beam displacer (BD) before entering the memory cell for storage and retrieval. Post-retrieval filtering comprises a Glan-Taylor polarizer (GTP), three double-passed etalons, and a band-pass filter (BP). The signal is finally routed to fiber beam-splitters and four superconducting nanowire single-photon detectors (SNSPDs) for detection. Employing multiple detectors helps mitigate detector dead time effects and enables higher overall count rates. (AWG: arbitrary waveform generator; PC: Pockels cell; PBS: polarizing beamsplitter; M: mirror; DM: dichroic mirror; HWP: half-wave plate; WPs: a half-wave plate and a quarter-wave plate.) (B) Pulse sequences. The diagram illustrates the timing of the signal and control pulses, along with the gating windows of the optical pumping and the Pockels cell.
@@
Our Raman memory setup is shown in Fig. 5(A). The storage medium consists of a 75 mm-long cesium (Cs-133) vapor cell with a three-layer $\mu$-metal magnetic shielding. We heated the cell to $105\pm 1$ °C using a quad-twisted cryogenic wire, achieving an effective optical depth of $(4.8\pm 0.1)\times 10^{3}$. The atoms were initialized to the initial state $|g\rangle$ by an external cavity diode laser (ECDL, Toptica DL100) driving the D1 transition ($6^{2}\textbf{S}_{1/2}\rightarrow 6^{2}\textbf{P}_{1/2}$). An electro-optical modulator (EOM 3, EOSpace 850UL) gated the pump laser, switching it off during the experimental window. We achieved a pumping efficiency of $(99.6\pm 0.1)$%.
@@
We operated our memory with a detuning $\Delta=2\Delta_{\text{hf}}$ to suppress four-wave-mixing noise [71]. Both the signal and control fields were derived from a single continuous wave (CW) ECDL laser (Toptica DL Pro) operating at the control frequency (351.7122 THz). An arbitrary waveform generator (AWG, Tektronix AWG70001A) drove the electro-optical modulator EOM 1 (Sacher Lasertechnik AM830PF) to carve the CW laser into pulses. The laser was then split using a 90:10 beam splitter, directing 90% of the optical power to the control field path. The remaining 10% went to EOM 2 (New Focus 4851), which generated a sideband at the signal frequency (351.7030 THz). We selected this sideband as the signal field by using a Fabry-Pérot etalon with a free spectral range (FSR) of 36.8 GHz. The signal field was delayed by 437 ns using a long fiber to achieve temporal synchronization with the control pulses, ensuring their temporal overlap for memory storage and retrieval. Before entering the memory, the signal field also passed through a Pockels cell to remove residual CW background in the retrieval window. The polarization of the signal and control fields was set to orthogonal for memory interaction, allowing their combination using a beam displacer before being focused into the vapor cell.
[…]
@@
After the memory, we suppressed the control and pump fields using a Glan-Taylor polarizer, followed by three double-passed etalons (two with FSR of 18.4 GHz and one with FSR 103 GHz) and a bandpass filter (central wavelength: 850 nm, FWHM Bandwidth: 10 nm). Finally, we split the signal into four equal parts and detected them using four superconducting nanowire single-photon detectors (SNSPDs, Photon Spot). Employing multiple detectors helps mitigate detector dead time effects and enables higher overall count rates. The detected counts were registered using a time tagger (Swabian TimeTagger20).
@@
$\Psi(t)=A\cos\left(\frac{\epsilon\sigma t-\phi}{2}\right)\exp(-t^{2}\sigma^{2}),$ (8)
@@
One key requirement for this experiment is the accurate generation of signal and control pulses. To ensure high-fidelity pulse carving, we characterized the frequency response of our electronic pulse generation system, which comprises the AWG, an RF splitter, and two RF amplifiers that drive the EOM 1. By measuring the system’s frequency response function, we applied a correction to the input signals of the AWG, compensating for frequency-dependent variations in the RF components. This correction ensures uniform amplification across all frequency components, minimizing distortions in the carved pulses. The measured intensity pulse shapes of HG0 and HG1 are presented in the Supplementary Information.
@@
[…]
@@ Acknowledgments
Supplementary Information: Super-resolving frequency measurement with mode-selective quantum memory
@@ IV.1 Fundamental bias at small separations
In the absence of crosstalk, the ideal detection probability for mode $n$ is
$P(n|\epsilon)=\frac{\epsilon^{2n}}{16^{n}n!}\exp\left(-\frac{\epsilon^{2}}{16}\right).$ (S1)
And the Fisher information for this HG measurement basis can be calculated as
$\mathcal{F}_{\text{HG}}(\epsilon)=\sum_{n=0}^{\infty}\frac{1}{P(n|\epsilon)}\left(\frac{\partial P(n|\epsilon)}{\partial\epsilon}\right)^{2}.$ (S2)
Considering only the first two modes, the normalized probability of projecting onto the HG1 mode after perturbation by the crosstalk matrix $\bm{M}$ is
$\tilde{P}(1|\epsilon)=\beta-\frac{16(\alpha+\beta-1)}{16+\epsilon^{2}},$ (S3)
and the HG0 projection probability is $\tilde{P}(0|\epsilon)=1-\tilde{P}(1|\epsilon)$. The exact expression for the Fisher information of the non-ideal two-mode projection method can be derived from Eq. S2 in the main text as follows:
@@
We assume the mode crosstalk between the two modes is low, i.e. $\alpha\approx 1$ and $\beta\approx 1$, and the separation is small. Then Eq. S4 can be approximated as Eq. 2 in the main text. Using maximum likelihood estimation (Eq. 3 in the main text), the likelihood function is given by
$\mathcal{L}=N_{0}\ln{\tilde{P}(0|\epsilon)}+N_{1}\ln{\tilde{P}(1|\epsilon)},$ (S5)
@@
Since we employ photon count detection, the detection counts $N_{0}$ and $N_{1}$ follow a Poisson distribution. Specifically, we model $N_{1}$ as $\text{Poisson}(\mu_{1})$, where $\mu_{1}=\tilde{P}(1|\epsilon)N$. Substituting this distribution into the estimator formula in Eq. S6, we derive the mean square (MSE) error as
@@
Figure S1: Simulation results of the mean square error (MSE) and the unbiased Cramér-Rao lower bound (CRLB) for maximum likelihood estimation. (A) Simulated MSE, where the white region indicates the onset of truncation due to the non-negativity constraint. (B) MSE with the non-negativity constraint applied. (C) The unbiased CRLB. (D) The difference between the CRLB and the MSE (CRLB - MSE).
The MSE simulated using our experimentally measured crosstalk is shown in Fig. S1 as a function of the separation $\epsilon$ and the number of photons $N$. In panel (A), the white region indicates where the MLE estimators begin to truncate negative estimates, and panel (B) displays the corresponding MSE with this truncation applied. Panel (C) presents the unbiased Cramér-Rao lower bound (CRLB). By taking the difference between the unbiased CRLB and the MSE (CRLB - MSE), we obtain panel (D), which highlights a transition: for large separations, the MSE exceeds the unbiased CRLB (blue regions), but as the separation decreases and with larger bias, it drops below the CRLB (yellow regions).
@@ IV.2 Quantum memory interaction
$N_{\text{in(out)}}=\int_{-\infty}^{\infty}d\tau\left<S_{\text{in(out)}}^{\dagger}(\tau)S_{\text{in(out)}}(\tau)\right$> (S10)
$N_{\text{stor}}=\int_{0}^{L}dz\left<B_{\text{stor}}^{\dagger}(z)B_{\text{stor}}(z)\right$>. (S11)
@@
The modal properties of the Raman memory can be understood by performing singular value decomposition on the Green functions $K_{1}$ and $K_{2}$ [55]. For the storage process, $K_{1}(z,\tau)$ can be decomposed into
@@ IV.4 Different retrieval modes and retrieval times
We also performed the experiment using different retrieval modes to demonstrate the flexibility in choosing the read-out control pulse. The results are shown in fig. S4, where the control read-out pulses are set to HG0 and HG1, respectively. The superresolution performance remains comparable in both cases, and the retrieved signals (the second red pulse) clearly exhibit the corresponding temporal mode shapes.
@@ IV.5 Sources of mode crosstalk
The primary source of temporal mode crosstalk arises from the memory’s storage interaction. Although the memory operates approximately in a single-mode regime, increasing the control field strength induces stronger AC Stark shifts. This leads to the coupling of orthogonal signal modes into the spin wave, distorting the ideal projection probabilities and introducing bias. Fig. S5 illustrates how different storage efficiencies (i.e. different control read-in pulse power) affect the superresolution performance and estimation precision relative to DI methods.
Panel (A) shows the raw estimators $\hat{\epsilon}=4\sqrt{N_{1}/N_{0}}$ for different storage efficiencies. As storage efficiency increases, so does crosstalk, resulting in greater bias in the raw estimators. In panel (B), we apply maximum likelihood estimation and plot the corresponding unbiased CRLB for the two-mode demultiplexing method, alongside the quantum CRLB and the CRLB for DI methods. Higher storage efficiency leads to an increase in the CRLB, resulting in reduced estimation precision. This trend is further quantified in panel (C), where the precision improvement factor, $\mathcal{F}/\mathcal{F}_{\text{DI}}$, is plotted as a function of separation $\epsilon$. As the storage efficiency decreases, the superresolution parameter increases nearly logarithmically. However, very high precision improvement using very low storage efficiencies is impractical in experiments, as fewer signal photons are retrieved, constant background and control field leakage reduce the signal-to-noise ratio, potentially increasing the observed crosstalk and estimation bias.